with the following information for a purchased part, what is the total annual cost for the item? (4 points) Annual demand = 200,000 units, Purchase price = $5 per unit, Order quantity = 10,000 units, Order cost = $8 per order, Annual holding cost = $3 per unit
Answer #1

Total number of orders = 200000/10000 = 20

Total ordering cost = 20*$8 =$160

Total product cost = 200000*$5= $1000000

Total holding cost = average inventory *$3

Average inventory = 10000/2 = 5000

=5000*$3

=$15000

Total annual cost of the item = $15000+$1000000+$160 = $1015160